Circular No. 7388

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 2000af AND 2000al A. V. Filippenko and A. L. Coil, University of California at Berkeley,

report that observations of two of the apparent supernovae reported by Schaefer on IAUC 7387 were obtained with the Shane 3-m reflector at Lick Observatory on Mar. 15. The data are not yet fully reduced. The spectrum of SN 2000af, taken in poor seeing, shows no obvious supernova features but does have a probable narrow [O II] 372.7-nm line for z = 0.13; galaxy light probably dominates this particular spectrum. The spectrum of SN 2000al has a deep absorption line at 737.0 nm, which gives z = 0.20 if taken to be the Si II line or z = 0.12 if taken to be the Hα line.

POSSIBLE NOVA IN NGC 3314 W. C. Keel, University of Alabama; and L. M. Frattare, Space Tele-

scope Science Institute, on behalf of the Hubble Heritage Team, report the detection of a possible nova in the overlapping galaxy pair NGC 3314A and 3314B. HST WFPC2 observations with the F450W, F555W, and F675W filters on Mar. 10.47–10.57 UT show a new stellar object that was not present on WFPC2 F450W (to B ∼ 25.5) and F814W images from 1999 Apr. 4. Standard (approximate) transformations to the UBV system give magnitudes B = 22.42, V = 21.64, R = 20.87. The new object is located at α = 10^{h}37^{m}12^{s}.82, δ = −27^{o}40′51′′.5 (equinox 2000.0), which is 1′′.2 west and 10′′.3 north of the optically brighter foreground nucleus of NGC 3314A (itself at position end figures 12^{s}.91, 41′01′′.8).

XTE J1859+226 P. Garnavich and J. Quinn, University of Notre Dame, report: “We

obtained 12.5 hr of photometry for this x-ray nova over five nights with the 1.8-m Vatican Advanced Technology Telescope beginning Mar. 16.4 UT. XTE J1859+226 showed variations of up to 0.3 mag over 2 hr and nightly variations of up to 0.5 mag in the V filter. Phasing of the variations suggests a period of 9.15 ± 0.05 hr; however, due to aliasing, periods of 14.79 and 6.60 hr are also possible. Our photometry rules out a current period of 6.734 hr reported in 1999 Nov. (IAUC 7303). The 0.5-mag rise between Mar. 16 and 17 suggests that long-term variations also are present or that XTE J1859+226 may be an eclipsing system.”

2000 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green